OpenStudy (123ab456c):

Which equation can be used to find the answer? Luna sells bracelets at the county fair. So far she has sold 23 bracelets and has 112 left. How many bracelets did she have before selling any? A. b ÷ 2 = 23 b = 46 46 bracelets B. b ÷ 2 = 112 b = 224 224 bracelets C. b + 23 = 112 b = 89 89 bracelets D. b – 23 = 112 b = 135 135 bracelets

OpenStudy (welshfella):

let b be the number of bracelets she had before she sold any. She sold 23 bracelets, therefore the number she has left is b - 23 - and this is equal to 112.

OpenStudy (123ab456c):

IS the answer C

OpenStudy (welshfella):

No - Look at my post again

OpenStudy (123ab456c):

the answer is B

OpenStudy (123ab456c):

NO- D

OpenStudy (123ab456c):

i think it is D

OpenStudy (welshfella):

I think you are guessing after 23 have been sold the number left is b take away 23 or b - 23. - and we are also told in the question that this number left is 112. so these 2 things are equal so b - 23 = what?

OpenStudy (123ab456c):

A

OpenStudy (123ab456c):

IS IT A

OpenStudy (123ab456c):

and im not gessing im trying to get help

OpenStudy (welshfella):

÷ means divide not subtract.

OpenStudy (welshfella):

And I'm trying to help you.

OpenStudy (123ab456c):

is the answer C

OpenStudy (welshfella):

no

OpenStudy (123ab456c):

i told you that the answer is D

OpenStudy (welshfella):

Sorry I missed your earlier post The answer is D b - 23 = 112 now add 23 to both sides b - 23 + 23 = 112 + 23 b = 135

OpenStudy (123ab456c):

its fine thanks for the help i told you the answer was D though

OpenStudy (welshfella):

yes I know - missed that.

OpenStudy (123ab456c):

lol mistakes happen

OpenStudy (welshfella):

they sure do